OpenStudy (anonymous):

substitution method for this proble, please help 7a=4-b 7a=6-b

OpenStudy (nubeer):

if 7a = 4-b and 7a =6-b 7a =7a 4-b = 6-b.. can u do it from here?

OpenStudy (anonymous):

2-b

OpenStudy (anonymous):

if that is not hte answer than no I cant this subject is not my best one

OpenStudy (nubeer):

hmm can u make sure the question again.. bcos if thats ur question b is cancelling out with is ntt possible.

jhonyy9 (jhonyy9):

what is the substitution method ?

jhonyy9 (jhonyy9):

@nubeer ATTENTION !!!

OpenStudy (anonymous):

I have the slightest idea please advise

OpenStudy (nubeer):

its a method in which we replace a variable with different expression.

OpenStudy (anonymous):

wondering if you guys could take me step by step on the last 14 question that I have for a final

OpenStudy (anonymous):

that is what I thought I was taking away the 7a and I was doing it like 6-4 but nothing is coming out right

jhonyy9 (jhonyy9):

yes @nubeer is right so calcule for example the x variable from the first equation and this replace in the second can you make this step by step for right,correct understanding of asker here now ?

OpenStudy (anonymous):

jhonyy9 could you help me out step by step not give me the answer but walk me through these last questions if possible

OpenStudy (anonymous):

?? that did not make since to me

OpenStudy (nubeer):

@bigmommatiad can you make sure agian that the question is written right because for your question the values are not coming..

OpenStudy (anonymous):

are you saying there is no solution tbecause that could be an answer

OpenStudy (anonymous):

that is the questions

OpenStudy (nubeer):

but i can tell u in general what is subsitution method. 7a = 4-b b = 4-7a in this step i have to make a or b the subject i have made b the subject.. now replace the "b" i found in other equation 7a = 6- (4-7a) the rest is arithmetic in this case "a" is cacelling out so i think no solution,.

OpenStudy (anonymous):

k dont leave have another question

OpenStudy (anonymous):

ok so I have x=7-y x y 0 ? 6 ? 7 ?

OpenStudy (nubeer):

hmm i see .. well why don't u try for me first.. write steps here if u do soemthign wrong i will correct u.. what u say?

OpenStudy (anonymous):

sounds good so for the first one would I do 7-0, then 7-6, then 7-7

OpenStudy (nubeer):

yup.. there u go.. u did it :)

OpenStudy (anonymous):

omg wow k hold on please

OpenStudy (anonymous):

the number of state and federal inmates in millions during year x, were x>2002 can be approx by following formula n=0.05x-98.65 determine year in which there were 1.5 million inmates

OpenStudy (anonymous):

i am completely lost on this one

OpenStudy (anonymous):

nubeer you still here

OpenStudy (nubeer):

hmm yes i am.. well look u have to find x.. so u have to rearrage your expression and make x the subject. n +98.65 =0.05x now divide both sides by 0.05 so u can get x as a subject.. do u get till this part?

OpenStudy (anonymous):

so 1973

OpenStudy (anonymous):

no that cant be right

OpenStudy (anonymous):

b=4-7a substitute this value in second equation 7a=6-(4-7a) 7a=6-4+7a 0=-2 thus not possible and hence no solution

OpenStudy (nubeer):

hmm no.. (n+98.65)/0.05 =x do u get it now?

OpenStudy (anonymous):

we are doing a different problen hr it is the number of state and federal inmates in millions during year x, were x>2002 can be approx by following formula n=0.05x-98.65 determine year in which there were 1.5 million inmates 6 minutes ago

OpenStudy (anonymous):

a little but I do not know what n is

OpenStudy (nubeer):

n=1.5.. as it already says n would be in millions..